OpenStudy (hotpinkgirl):

find the funtion rule x = -2, -1, 0, 1, 2, y= 9, 4, -1, -6, -11

OpenStudy (hotpinkgirl):

@iambatman @shrutipande9 @perl @sleepyjess please help!!!!!!!!!!!!!!!

OpenStudy (hotpinkgirl):

@KyanTheDoodle @NeonStrawsForever

OpenStudy (kyanthedoodle):

I think I may be able to help

OpenStudy (hotpinkgirl):

ok, thanks

OpenStudy (kyanthedoodle):

What patterns can you find by this rule?

OpenStudy (kyanthedoodle):

I found that as x goes up by 1, y goes down by 5.

OpenStudy (hotpinkgirl):

ya thats right

OpenStudy (hotpinkgirl):

so what would be the funtion rule?

OpenStudy (kyanthedoodle):

That's your question to answer. What's the function rule?

OpenStudy (hotpinkgirl):

i think the rule is y= 5??

OpenStudy (kyanthedoodle):

I don't think that's it. What kind of answer are they looking for?

OpenStudy (hotpinkgirl):

you have to find the funtion rule and explain how you did it

OpenStudy (kyanthedoodle):

Well I don't really know how to explain how I found it. You just kinda see it at one point. Any more ideas on what the function rule is. It goes along the lines of y = x (and then something)

OpenStudy (hotpinkgirl):

i dont know

OpenStudy (kyanthedoodle):

Hmm

OpenStudy (hotpinkgirl):

thank i need to go

OpenStudy (kyanthedoodle):

x = 1, y = -5 because as I said before, as X goes 1 up, Y goes 5 down.

OpenStudy (amistre64):

notice that for values of x: y(0) = -1 -5(0) y(1) = -1 -5(1) y(2) = -1 -5(2) etc .... therefore, what does y(x) equal?

OpenStudy (nincompoop):

find the common difference :)

OpenStudy (anonymous):

Here is an idea:\[ y=f(x) =\begin{cases} -1&x=0\\ f(x-1)-5&\text{otherwise} \end{cases} \]